OpenStudy (anonymous):

(sqrt)- 6(2 + (sqrt) - 6))

OpenStudy (anonymous):

\[\sqrt{-6}\left( 2 +\sqrt{-8} \right)\]

OpenStudy (anonymous):

What format is the answer supposed to be in? Does it want a decimal, or fraction, or what?

OpenStudy (anonymous):

just solve

OpenStudy (anonymous):

No instructions at all? Cause if you just solve it'll come out to be one lonnnnnnggggg decimal.

OpenStudy (anonymous):

But if you keep it in square root form it'll be cleaner. It just depends on what your teacher is looking for.

OpenStudy (anonymous):

-4(sqrt) 3 + 21(sqrt)6 is one of the answers

OpenStudy (anonymous):

choices*

OpenStudy (anonymous):

Okay so it wants it just like that alright. Well you have your equation. \[\sqrt{-6}(2+\sqrt{-8})\] You need to multiply everything by \[\sqrt{-6}\] So it'll look like this,\[(\sqrt{-6}\times2)+(\sqrt{-6}\times \sqrt{-8})\]

OpenStudy (anonymous):

Can you take it one step further from there?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Alright well \[2\times \sqrt{-6}\] can't be simplified any further. So you have 1/2 of your answer.

OpenStudy (anonymous):

Now for the second part of your answer,\[\sqrt{-6} \times \sqrt{-8}\] To multiply square roots you multiply the numbers within their sign. soooo\[\sqrt{-6(-8)}\] Now you solve that part.